What is the effective annual rate for an APR of 15.70 percent compounded monthly?
Answer #1

APR = 15.70%p.a. (given)

Compounding = monthly

i/r = 15.70% / 12 months = 1.308333

EAR (Effective Annual rate) = (1+i/r)^n - 1

= (1+1.30833%)^12 - 1

= 16.88%

Hence, EAR for an APR of 15.7% compounded monthly is equal to 16.88%